1. A particle of mass m moves according to x = x0 + at2, y = bt3, and z = ct. Verify that


                                      dLdt=r×F=τ                      


1
Expert's answer
2021-01-11T11:42:06-0500

Answer

Position vector can be written as

"r=x\hat {i}+y\hat {i}+z\hat {i}"

="(x_0 + at^2) \hat {i} +bt^3 \hat {j}+ct
 \hat {k}"

Now acceleration can be written as

"a=\frac{d^2r}{dt^2}= (2a) \hat {i} +6bt\hat {j}"

Now cross product

"r\u00d7F=r\u00d7(ma)"

"=m(r\u00d7a)"

"=\begin{vmatrix}
 \hat{i} & \hat{j} & \hat{k}& \\
 (x_0 + at^2) & bt^3 & ct \\
 2a & 6bt &0
\end{vmatrix}"

"=-6bct^2\hat{i}-2act \hat{j}+(6x_0bt+4abt^3) \hat{k}"

This means

"r\u00d7F \u2260 0"

In this specific case we can rewrite as

"dL\/dt=r\u00d7F=\u03c4"